# Generalist Bimanual Manipulation via Foundation Video Diffusion Models

### v1

@@ Abstract
Bimanual robotic manipulation, which involves the coordinated control of two robotic arms, is foundational for solving challenging tasks. Despite recent progress in general-purpose manipulation, data scarcity and embodiment heterogeneity remain serious obstacles to further scaling up in bimanual settings. In this paper, we introduce VIdeo Diffusion for Action Reasoning (VIDAR), a two-stage framework that leverages large-scale, diffusion-based video pre-training and a novel masked inverse dynamics model for action prediction. We pre-train the video diffusion model on 750K multi-view videos from three real-world bimanual robot platforms, utilizing a unified observation space that encodes robot, camera, task, and scene contexts. Our masked inverse dynamics model learns masks to extract action-relevant information from generated trajectories without requiring pixel-level labels, and the masks can effectively generalize to unseen backgrounds. Our experiments demonstrate that with only 20 minutes of human demonstrations on an unseen robot platform (only 1% of typical data requirements), VIDAR generalizes to unseen tasks and backgrounds with strong semantic understanding, surpassing state-of-the-art methods. Our findings highlight the potential of video foundation models, coupled with masked action prediction, to enable scalable and generalizable robotic manipulation in diverse real-world settings.
Figure 1: Overview of VIdeo Diffusion for Action Reasoning (VIDAR), a foundation video diffusion model for video prediction and masked inverse dynamics model (MIDM) for action regression. Trained on 750K multi-view bimanual videos with test-time scaling (TTS) applied during testing, VIDAR can adapt to new robot platforms with 20 minutes of demonstrations with state-of-the-art performance and generalize to unseen tasks with strong semantic understanding.
@@ 1 Introduction
[…]
To address these limitations, we propose VIdeo Diffusion for Action Reasoning (VIDAR), as is shown in Figure 1. We train a foundation video diffusion model with 750K multi-view bimanual videos collected from three real-world robot platforms, where embodiment heterogeneity is handled by a unified observation space. Our key insight is that pre-trained video prediction models can serve as transferable priors for robotic control, provided they are adapted to the dynamics of physical interaction and grounded in action-relevant content. We further adopt test-time scaling [openai-o1] to enhance the performance of the foundation model when domain-specific data is scarce.
With limited human demonstrations, it is also non-trivial to translate domain-specific videos into corresponding actions. To solve this problem, we introduce a Masked Inverse Dynamics Model (MIDM) that learns to focus on and extract action-relevant regions in the generated frames, without requiring any pixel-level supervision. At test time, the learned masks generalize well in discarding action-irrelevant features, further enhancing stability and generalization.
[…]
@@
  • We present the first large-scale video foundation model for bimanual manipulation, trained on 750K diverse demonstrations across three robot platforms with a unified observation space.
@@ 2 Method
Figure 2: We introduce a unified observation space that integrates multi-view robotic videos and language instructions. This enables the training of a foundation video diffusion model on 750K multi-view bimanual robotic episodes, allowing it to generalize to unseen robot platforms with only 20 minutes of human demonstrations. At inference time, it generates multiple candidate videos, then applies test-time scaling (TTS) to select the most suitable one. We also propose the masked inverse dynamics model (MIDM), which learns masks to identify and attend to action-relevant regions for robust frame-based action regression.
In this section, we formally define our problem and describe the proposed VIDAR approach in detail. The overview of our method is shown in Figure 2.
@@ 2.1 Problem Formulation
Let $\mathcal{L}$ denote the language instruction space, $\mathcal{O}$ the observation space, and $\mathcal{A}$ the action space. Our objective is to learn a conditional robot manipulation policy $\pi:\mathcal{L}\times\mathcal{O}\rightarrow\mathbb{P}(\mathcal{A})$, where $\mathbb{P}(\cdot)$ represents the set of all probability measures over the given space. Directly learning this policy poses significant challenges, mainly due to the large volume of demonstrations required to achieve generation and generalization over three different modalities: language, observation, and action. For example, RDT-1B [rdt-1b] collects 27 hours of human demonstrations for adaptation to the Aloha robot, which greatly limits its applicability to new platforms.
@@
We choose the Aloha platform with three views as our bimanual robot in the target domain, since it is widely used and suitable for teleoperations [DBLP:conf/rss/ZhaoKLF23, rdt-1b]. For our target robot, actions are defined as target joint/gripper positions without conditioning on the history, which means the inverse dynamics model can be viewed as sequential applications of a simpler mapping from one image to one action. With a little abuse of notation, we denote the simpler mapping as the inverse dynamics model throughout.
@@ 2.2 Video Generation Model
We adopt rectified flow [lipman2022flow, liu2022flow] models, which generate high-quality videos by modeling pixel flow over time. Specifically, the model parameterizes a flow function $v:\mathcal{V}\times\mathbb{R}\times\mathcal{L}\rightarrow\mathcal{V}$, which models the velocity of pixels as they transition from noisy video frames to target frames under certain conditions and time, leading to the following ODE over a video $x_{t}$ (with $x_{0}$ sampled from a Gaussian and $x_{1}$ as the output video):
@@
To address the diversity and heterogeneity of real-world bimanual robots, we pre-train our video generation model on a large-scale corpus of approximately 750K open-sourced episodes, all projected into a unified observation space. This dataset includes episodes with three fixed camera views, a configuration common in bimanual setups [mobile_aloha, agibot]. The primary challenge lies in harmonizing heterogeneous embodiments: different robots vary in camera placements, viewing angles, and sensor resolutions, leading to covariate shifts in perception. Moreover, wrist-mounted views often introduce task-irrelevant information, increasing input noise and model variance.
@@ 3.1.1 Datasets
For pre-training, we utilize episodes sourced from Agibot-World [agibot], RoboMind [robomind], and RDT [rdt-1b]. For Agibot-World, episodes are segmented into shorter clips using their frame-level annotations. As is required by the unified observation space, we provide descriptions of the robot type and camera placements in addition to the provided task instructions. The resulting pre-training dataset comprises 746,533 episodes.
@@ 3.1.2 Training and Inference
We utilize Vidu 2.0 [bao2024viduhighlyconsistentdynamic] as our video generation model. We train for 23,000 iterations with a batch size of 128, where the first 10,000 steps involve pre-training on the diverse robotics dataset, followed by 13,000 fine-tuning steps. To reduce inference costs, we uniformly downsample the training videos to 8 frames per second (fps). For the masked inverse dynamics model, we adopt the U-Net [unet] structure as the mask prediction network and the ResNet [resnet] structure as the action regression network, with $\lambda=3\times 10^{-3}$ (effects of $\lambda$ are shown in Appendix C). We train it exclusively on the fine-tuning dataset. We use AdamW [adamw] for all our training.
We use open-loop control for VIDAR; the videos are generated in a single batch, without subsequent generation after the initial run. Using 8 NVIDIA Ampere-series 80GB GPUs, generating one video with 60 frames (7.5 seconds duration at 8fps) costs about 25 seconds. For test-time scaling, we choose $K=3$, and three videos with different random seeds are generated in parallel, evaluated by GPT-4o [gpt-4o]. More details of training and inference can be found in Appendix B.
We acknowledge that Vidu 2.0 is not open source. To demonstrate the generality of our approach, we also perform small-scale experiments using the open-source HunyuanVideo model [hunyuanvideo]. Related results are shown in Appendix E.
@@ 3.1.3 Baselines
We use the same checkpoint for the video generation model as our method, and train a diffusion model to generate short action sequences based on the latent features during one-step video generation and CLIP [clip] embeddings of the task instructions. Notably, they use closed-loop control, which means new action sequences are generated and executed after previous executions.
@@ 3.2 Experimental Results
We test two baseline methods and our method over 16 tasks under three scenarios: seen task and background (six tasks), unseen task (four tasks), and unseen background (six tasks). Detailed explanations are as follows:
@@
We evaluate the video generation quality of the original Vidu 2.0 model and our pre-trained version over the unseen target domain using VBench [vbench]. As is shown in Table 2, we find that pre-training using large-scale robotic videos under our unified observation space enhances both the consistency and quality of generated frames, which are important for robot control tasks.
@@ 4 Related Work
Vision-Language-Action (VLA) models integrate perception, language understanding, and action generation to enable general-purpose embodied intelligence. However, existing VLA systems rely heavily on large-scale, task-specific datasets—often requiring hundreds of thousands of trajectories—to achieve robust performance. For instance, RT-1 [rt1] uses 130K real-world episodes, while RT-2 [rt2] scales to 1B image-text pairs. Recent works like OpenVLA [openvla], Octo [octo], Pi0 [pi0], and RDT-1B [rdt-1b] further expand to millions of demonstrations across diverse embodiments. Despite these efforts, such data-intensive pipelines present a scalability bottleneck and limit generalization to novel tasks and domains, motivating more data-efficient and transferable alternatives.
@@
Video generation models for embodied AI predict future scene dynamics to assist robotic planning and policy learning. Current works such as UniPi [unipi], RoboDreamer [robodreamer], Gen2Act [gen2act], and GR-1 [gr1] mainly adopt text-conditioned video generation with a fixed single-camera view, demonstrating how a single-arm robot physically compliantly executes a series of actions. Additionally, Genie 2 [genie2] generates interactive 3D environments via autoregressive latent diffusion, enabling scalable training for embodied agents. These models implicitly encode physical laws through video synthesis, reducing reliance on real-world robotics data. However, current works mostly remain limited to single-arm robots and rely on a single camera view, which restricts their applicability in more complex real-world scenarios.
@@
Recent efforts couple video generation with action synthesis to enhance physical consistency and interoperability. For instance, Video-Prediction-Policy (VPP) [vpp] learns implicit inverse dynamics conditioned on future representations predicted by video diffusion model, while UVA [uva] unifies video-action latent spaces with lightweight diffusion heads. Other works, such as VidMan [vidman] and Dreamitate [dreamitate], integrate low-level action representations with video prediction to predict actions. These approaches bridge spatiotemporal dynamics between vision and action, offering novel solutions for complex manipulation tasks. However, these approaches require end-to-end joint training of video generation and action prediction models, which limits their flexibility and adaptability. Integrating a more advanced video generation model would necessitate retraining the entire system.
@@ Appendix E Reproduction using the HunyuanVideo Model
AdamW Weight Decay 00
@@ Appendix F Hardware Details
Degree of Freedom $2\times(6+1)=14$

### v2

# Vidar: Embodied Video Diffusion Model for Generalist Bimanual Manipulation

@@ Abstract
Bimanual robotic manipulation, which involves the coordinated control of two robotic arms, is foundational for solving challenging tasks. Despite recent progress in general-purpose manipulation, data scarcity and embodiment heterogeneity remain serious obstacles to further scaling up in bimanual settings. In this paper, we present Video Diffusion for Action Reasoning (Vidar), a two-stage framework that leverages large-scale, diffusion-based video pre-training and a novel masked inverse dynamics model for action prediction. We pre-train a general video diffusion model on 750K multi-view videos from three real-world bimanual robot platforms, utilizing a unified observation space that encodes robot, camera, task, and scene contexts. Our masked inverse dynamics model learns masks to extract action-relevant information from generated trajectories without requiring pixel-level labels, and the masks can effectively generalize to unseen backgrounds. Experimental results demonstrate that with only 20 minutes of human demonstrations on an unseen robot platform (only 1% of typical data requirements), Vidar generalizes to unseen tasks and backgrounds with strong semantic understanding, surpassing state-of-the-art methods. Our findings highlight the potential of video foundation models, coupled with masked action prediction, to enable scalable and generalizable robotic manipulation in diverse real-world settings.
Figure 1: Overview of Video Diffusion for Action Reasoning (Vidar), a video diffusion foundation model for video prediction and masked inverse dynamics model (MIDM) for action regression. Trained on 750K multi-view bimanual videos with test-time scaling (TTS) applied during testing, Vidar can adapt to new robot platforms with 20 minutes of demonstrations with state-of-the-art performance and generalize to unseen tasks with strong semantic understanding.
@@ 1 Introduction
Bimanual manipulation—the coordinated control of two robotic arms—is central to many high-precision tasks in both industrial automation and domestic robotics. In contrast to single-arm manipulation, dual-arm skills enable stable object handling, in-hand reorientation, and multi-point contact control, which are essential for tasks such as cloth folding, liquid pouring, or tool-assisted assembly. Recent work on vision-language-action (VLA) foundation models [1, 2, 3, 4] has demonstrated promising results by leveraging large-scale multimodal pre-training, but extending these models to general-purpose bimanual settings remains highly nontrivial.
The challenge lies in the inherent complexity of bimanual control: the action space scales combinatorially with the number of controllable joints, while task success demands tight temporal coordination between arms, precise contact dynamics, and long-horizon planning. These properties substantially increase the data requirements and sensitivity to platform-specific variations. In practice, progress is limited by two fundamental bottlenecks. First, data scarcity: collecting large-scale, high-quality bimanual demonstrations is labor-intensive, cost-expensive and hardware-specific, resulting in orders-of-magnitude less data than what is available for single-arm or vision-language domains. Second, embodiment heterogeneity: differences in robot morphology, control frequency and kinematics typically lead to poor cross-platform transferability, requiring extensive robot-specific fine-tuning to adapt pre-trained models [5, 6].
To address these limitations, we propose Video Diffusion for Action Reasoning (Vidar) to effectively leverage Internet-scale video data for generalist bimanual manipulation. As outlined in Figure 1, Vidar consists of two ky stages. It first re-trains a general video diffusion model with 750K multi-view bimanual videos collected from three real-world robot platforms, where embodiment heterogeneity is handled by a unified observation space. We further adopt test-time scaling [7] to enhance the performance of the foundation model when domain-specific data is scarce. The predicted videos are then decoded into actions by the Masked Inverse Dynamics Model (MIDM), which learns to focus on and robustly extract action-relevant regions in the generated frames, without requiring any pixel-level supervision.
Such two-stage design makes Vidar be able to fully explore the embodied data pyramid, which consists of three progressively structured levels: (1) Web Video Data, offering broad, unstructured task diversity from Internet-scale sources; (2) Heterogeneous Robotic Data, providing mid-level embodied videos across diverse robot morphologies; and (3) Embodied Human Demonstrations, delivering fine-grained, domain-specific examples. In particular, the web video data is used for vanilla pretraining of the video diffusion model, endowing the model with general semantic understanding and visual rendering capabilities; heterogeneous robotic data supports embodied pretraining, enhancing its grasp of robot-specific dynamics and manipulation tasks; and embodied human demonstrations are used for fine-tuning, allowing precise adaptation to the target domain. By harnessing the first two levels of the pyramid, Vidar achieves strong generalization with significantly greater domain data efficiency—often by orders of magnitude.
Empirically, Vidar achieves strong performance with only 20 minutes of human demonstrations (3 demonstrations per task on average) in a previously unseen robotic platform. Notably, the amount of embodied human demonstrations required by Vidar is remarkably low—only 1/81 of that used by RDT [3] (27 hours), and 1/1200 of that used by $\pi_{0.5}$ [8] (400 hours). Moreover, Vidar outperforms state-of-the-art baselines VPP [9] and UniPi [10] by 58% and 40% respectively. Also, it can generalize to novel tasks and backgrounds, with good semantic understanding.
@@
  • We present the first large-scale embodied video diffusion model for bimanual manipulation, trained on 750K diverse demonstrations across three robot platforms with a unified observation space.
@@ 2 Method
Figure 2: We introduce a unified observation space that integrates multi-view robotic videos and language instructions. This enables the training of a video diffusion foundation model on 750K multi-view bimanual robotic episodes, allowing it to generalize to unseen robot platforms with only 20 minutes of human demonstrations. At inference time, it generates multiple candidate videos, then applies test-time scaling (TTS) to select the most suitable one. We also propose the masked inverse dynamics model (MIDM), which learns masks to identify and attend to action-relevant regions for robust frame-based action regression.
In this section, we formally define our problem and present Vidar in detail. The overview of our method is shown in Figure 2.
@@ 2.1 Problem Formulation
Let $\mathcal{L}$ denote the language instruction space, $\mathcal{O}$ the observation space, and $\mathcal{A}$ the action space. Our objective is to learn a conditional robot manipulation policy $\pi:\mathcal{L}\times\mathcal{O}\rightarrow\mathbb{P}(\mathcal{A})$, where $\mathbb{P}(\cdot)$ represents the set of all probability measures over the given space. Directly learning this policy poses significant challenges, mainly due to the large volume of demonstrations required to achieve generation and generalization over three different modalities: language, observation, and action. For example, RDT-1B [3] collects 27 hours of human demonstrations for adaptation to the Aloha robot, which greatly limits its applicability to new platforms.
@@
We choose the Aloha platform with three views as our bimanual robot in the target domain, since it is widely used and suitable for teleoperations [11, 3]. For our target robot, actions are defined as target joint/gripper positions without conditioning on the history, which means the inverse dynamics model can be viewed as sequential applications of a simpler mapping from one image to one action. With a little abuse of notation, we denote the simpler mapping as the inverse dynamics model throughout.
@@ 2.2 Video Generation Model
We adopt rectified flow [12, 13] models, which generate high-quality videos by modeling pixel flow over time. Specifically, the model parameterizes a flow function $v:\mathcal{V}\times\mathbb{R}\times\mathcal{L}\rightarrow\mathcal{V}$, which models the velocity of pixels as they transition from noisy video frames to target frames under certain conditions and time, leading to the following ODE over a video $x_{t}$ (with $x_{0}$ sampled from a Gaussian and $x_{1}$ as the output video):
@@
To address the diversity and heterogeneity of real-world bimanual robots, we pre-train our video generation model on a large-scale corpus of approximately 750K open-sourced episodes, all projected into a unified observation space. This dataset includes episodes with three fixed camera views, a configuration common in bimanual setups [14, 15]. The primary challenge lies in harmonizing heterogeneous embodiments: different robots vary in camera placements, viewing angles, and sensor resolutions, leading to covariate shifts in perception. Moreover, wrist-mounted views often introduce task-irrelevant information, increasing input noise and model variance.
@@ 3.1.1 Datasets
For pre-training, we utilize episodes sourced from Agibot-World [15], RoboMind [5], and RDT [3]. For Agibot-World, episodes are segmented into shorter clips using their frame-level annotations. As is required by the unified observation space, we provide descriptions of the robot type and camera placements in addition to the provided task instructions. The resulting pre-training dataset comprises 746,533 episodes.
@@ 3.1.2 Training and Inference
We utilize Vidu 2.0 [16] as our video generation model. We train for 23,000 iterations with a batch size of 128, where the first 10,000 steps involve pre-training on the diverse robotics dataset, followed by 13,000 fine-tuning steps. To reduce inference costs, we uniformly downsample the training videos to 8 frames per second (fps). For the masked inverse dynamics model, we adopt the U-Net [17] structure as the mask prediction network and the ResNet [18] structure as the action regression network, with $\lambda=3\times 10^{-3}$ (effects of $\lambda$ are shown in Appendix C). We train it exclusively on the fine-tuning dataset. We use AdamW [19] for all our training.
We use open-loop control for Vidar; the videos are generated in a single batch, without subsequent generation after the initial run. Using 8 NVIDIA Ampere-series 80GB GPUs, generating one video with 60 frames (7.5 seconds duration at 8fps) costs about 25 seconds. For test-time scaling, we choose $K=3$, and three videos with different random seeds are generated in parallel, evaluated by GPT-4o [20]. More details of training and inference can be found in Appendix B.
We acknowledge that Vidu 2.0 is not open source. To demonstrate the generality of our approach, we also perform small-scale experiments using the open-source HunyuanVideo model [21]. Related results are shown in Appendix E.
@@ 3.1.3 Baselines
We use the same checkpoint for the video generation model as our method, and train a diffusion model to generate short action sequences based on the latent features during one-step video generation and CLIP [22] embeddings of the task instructions. Notably, they use closed-loop control, which means new action sequences are generated and executed after previous executions.
@@ 3.2 Experimental Results
We test two baseline methods and our method under three scenarios: seen task and background (six tasks), unseen task (five tasks), and unseen background (six tasks). Detailed explanations are as follows:
@@
We evaluate the video generation quality of the original Vidu 2.0 model and our pre-trained version over the unseen target domain using VBench [23]. As is shown in Table 2, we find that pre-training using large-scale robotic videos under our unified observation space enhances both the consistency and quality of generated frames, which are important for robot control tasks.
@@ 4 Related Work
Vision-Language-Action (VLA) models integrate perception, language understanding, and action generation to enable general-purpose embodied intelligence. However, existing VLA systems rely heavily on large-scale, task-specific datasets—often requiring hundreds of thousands of trajectories—to achieve robust performance. For instance, RT-1 [24] uses 130K real-world episodes, while RT-2 [25] scales to 1B image-text pairs. Recent works like OpenVLA [2], Octo [26], Pi0 [27], and RDT-1B [3] further expand to millions of demonstrations across diverse embodiments. Despite these efforts, such data-intensive pipelines present a scalability bottleneck and limit generalization to novel tasks and domains, motivating more data-efficient and transferable alternatives.
@@
Video generation models for embodied AI predict future scene dynamics to assist robotic planning and policy learning. Current works such as UniPi [28], RoboDreamer [29], Gen2Act [30], and GR-1 [31] mainly adopt text-conditioned video generation with a fixed single-camera view, demonstrating how a single-arm robot physically compliantly executes a series of actions. Additionally, Genie 2 [32] generates interactive 3D environments via autoregressive latent diffusion, enabling scalable training for embodied agents. These models implicitly encode physical laws through video synthesis, reducing reliance on real-world robotics data. However, current works mostly remain limited to single-arm robots and rely on a single camera view, which restricts their applicability in more complex real-world scenarios.
@@
Recent efforts couple video generation with action synthesis to enhance physical consistency and interoperability. For instance, Video-Prediction-Policy (VPP) [9] learns implicit inverse dynamics conditioned on future representations predicted by video diffusion model, while UVA [33] unifies video-action latent spaces with lightweight diffusion heads. Other works, such as VidMan [34] and Dreamitate [35], integrate low-level action representations with video prediction to predict actions. These approaches bridge spatiotemporal dynamics between vision and action, offering novel solutions for complex manipulation tasks. However, these approaches require end-to-end joint training of video generation and action prediction models, which limits their flexibility and adaptability. Integrating a more advanced video generation model would necessitate retraining the entire system.
@@ Appendix E Reproduction using the HunyuanVideo Model
AdamW Weight Decay 0
@@ Appendix F Hardware Details
Degree of Freedom 2×(6+1)=142$\times$(6+1)=142 × ( 6 + 1 ) = 14